# Web Artifact Attacks Disrupt Vision Language Models

### v1

@@ Abstract
Vision-language models (VLMs) (e.g., CLIP, LLaVA) are trained on large-scale, lightly curated web datasets, leading them to learn unintended correlations between semantic concepts and unrelated visual signals. These associations degrade model accuracy by causing predictions to rely on incidental patterns rather than genuine visual understanding. Prior work has weaponized these correlations as an attack vector to manipulate model predictions, such as inserting a deceiving class text onto the image in a “typographic” attack. These attacks succeed due to VLMs’ text-heavy bias—a result of captions that echo visible words rather than describing content. However, this attack has focused solely on text that matches the target class exactly, overlooking a broader range of correlations, including non-matching text and graphical symbols, which arise from the abundance of branding content in web-scale data. To address this gap, we introduce “artifact-based” attacks: a novel class of manipulations that mislead models using both non-matching text and graphical elements. Unlike typographic attacks, these artifacts are not predefined, making them harder to defend against but also more challenging to find. We address this by framing artifact attacks as a search problem and demonstrate their effectiveness across five datasets, with some artifacts reinforcing each other to reach 100% attack success rates. These attacks transfer across models with up to 90% effectiveness, making it possible to attack unseen models. To defend against these attacks, we extend prior work’s artifact aware prompting to the graphical setting. We see a moderate reduction of success rates of up to 15% relative to standard prompts, suggesting a promising direction for enhancing model robustness.
@@ 1 Introduction
Figure 1: We propose a novel class of artifact-based attacks that expand the attack surface of Vision-Language Models beyond prior typographic attacks. For a sample task of predicting a person’s age, (a) illustrates typographic attacks from prior work [2] which manipulate predictions using class-matching text. In contrast, our proposed web artifact attacks demonstrate that non-matching text (b), standalone graphics (c), and graphics with embedded text (d) can also mislead the model.
@@
Vision-language models (VLMs), such as CLIP [26] and LLaVA [18], are trained on large-scale, lightly curated web datasets [29, 31]. Since web captions often describe what is present rather than what is important, models learn spurious patterns—unintended correlations that arise from frequent but unrelated co-occurrences [36, 1, 11, 13, 33, 32, 22, 15, 17]. For example, prior work found that Chinese characters frequently co-occur with the word “carton” in ImageNet [15], leading the model to misclassify images containing Chinese characters as a "carton," regardless of its actual content. Moreover, these correlations can be weaponized by attackers to manipulate model predictions. Fig. 1(a) shows a well-documented example, typographic attacks [25, 2], where inserting the text of the wrong category into an image tricks the model into predicting that category. This attack takes advantage of the model’s over-reliance on text, a bias reinforced by pretraining datasets in which captions frequently parrot text visible in the images [17].
@@
However, prior work has focused only on explicit class-matching text, overlooking a broader set of vulnerabilities, including non-matching text and graphical artifacts such as logos. These elements, which are common in web-scale datasets due to branding and advertisements, may introduce unintended visual cues that models may rely on for classification. To address this gap, we introduce artifact-based attacks, a broader class of attacks that exploit both non-matching text (Fig. 1(b)) and graphics (Fig. 1(c) and (d)) to mislead VLM predictions. To better understand their effects, we categorize graphical artifacts into two types: those with embedded text, which may reinforce textual biases, and those without text, isolating the impact of purely visual cues.
@@
Unlike typographic attacks [25, 2], where the adversary inserts class-matching text to deliberately exploit the model’s reliance on words, artifact-based attacks pose a greater threat because the defender does not know in advance the attacking artifacts. Thus, it is not obvious how to identify effective attacking artifacts or defend against them. To find attacks, we frame artifact-based attacks as a search problem, scanning image-text datasets [28, 5] to uncover text and graphics that can distort model predictions. The simplicity of these attacks makes them a notable security concern for deployed Vision-Language Models, as they only require dataset queries using similarity retrieval, making them highly accessible.
@@
To that end, our search process consists of three stages: (1) artifact retrieval, where we extract candidate artifacts by scanning image-text datasets for images with text or graphics (Fig. 2 Green) (2) effect estimation, where we apply artifacts to images from opposing classes and evaluate their influence on model predictions (Fig. 2 Purple), and (3) placement optimization, where we refine the positioning of artifacts to maximize attack success.
@@
We evaluate the resulting attack artifacts across different VLM training objectives, visual encoder architectures, and pretraining dataset curation. Our findings demonstrate that artifact-based attacks are highly effective across various vision-language model training paradigms, including contrastive learning [26], sigmoid loss [35], and large language model fusion with visual features [18, 37]. Furthermore, these vulnerabilities persist regardless of the visual encoder size (ViT-B-32, ViT-B-16, ViT-L-14), and attempts at stronger pretraining dataset curation (DataComp [10]).
@@
Finally, we explore mitigation strategies by building on prior work [7], which showed that incorporating typographic attacks into the prompt can reduce their effectiveness. For example, if an image of a youth is attacked with the text "senior," the prompt "an image of a senior with the word ’youth’ written on it" can weaken the attack. However, unlike typographic attacks, our attacks include graphical elements that cannot be easily incorporated into text. To address this, we propose augmenting the prompt with descriptions of graphical artifacts obtained through a captioning model. For LLM-based VLMs, we prompt the model to identify and describe these artifacts in its response. This approach reduces success rates by up to 15% relative to standard prompts, suggesting a promising direction for enhancing model robustness.
@@
  • We identify artifact-based attacks, which extend beyond typographic attacks by leveraging non-matching text, standalone graphics, and graphics with embedded text to mislead VLM predictions.
@@ 2 Related Work
Figure 2: Overview of the Attack’s Artifacts Search and Selection. The process involves two steps: Step 1: We filter an image-text dataset using CLIP-based retrieval and a set of artifact prompts. Retrieved artifacts are categorized into three groups: graphics with text ($\mathcal{A}^{\text{graphics-text}}$), graphics without text ($\mathcal{A}^{\text{graphics-no-text}}$), and unrelated text ($\mathcal{A}^{\text{text}}$) using OCR-based text extraction (Sec. 3.1.1). Step 2: artifacts are applied to opposing-class images (e.g., "Smiling" images when targeting "Frowning"), and their effectiveness is evaluated based on their ability to mislead model predictions. Artifacts are ranked according to their attack success rate, resulting in a refined set of high-impact artifacts for each category ($\mathcal{A}_{c^{*}}^{\text{graphics-text}}$, $\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$, and $\mathcal{A}_{c^{*}}^{\text{text}}$) (Sec. 3.1.2)
Relation to Adversarial Attacks. Adversarial attacks [34, 27, 38, 9, 6] learn an imperceptible noise, that disrupts the model visual recognition capabilities when added to the model. More recently, adversarial techniques have been weaponized for jailbreaking multimodal models to bypass safety mechanisms and generating restricted outputs [23, 30, 3]. While our attacks share a similar end goal to adversarial attacks, namely, disrupting model behavior, adversarial and artifact attacks are fundamentally different from each other. Indeed, our attacks originate as a result of spurious correlation in uncurated pretraining datasets, while adversarial attacks result from an artificial noise uniquely designed through the use of model intrinsics (e.g., gradients) to disrupt the fundamental mechanics of neural nets. Therefore, to develop such attacks, an attacker needs a sophisticated understanding of the intrinsic mechanics of neural nets such as e.g., Gradient [34, 27, 38], Meta Learning [9] or Attention maps [6] to name a few. However, the attacker, in our setup, only needs to know how to query the model.
@@ 3 Web Artifact Attacks Against Vision-Language Models
We propose artifact-based attacks, a new vector of unpredictable class of attacks that leverage artifacts from image-text datasets [28, 5]. Unlike typographic attacks, these artifacts—such as non-matching text (text that contains partial overlaps, phonetic similarities, or visually similar characters but does not explicitly spell out the target class), graphical symbols—do not need to explicitly match the target class yet can still mislead model predictions.
@@
Because these artifacts are not known a priori, they are harder to defend against while also making it challenging to determine in advance which artifacts will be effective. Therefore, we frame finding our attack artifacts as a search problem: systematically identifying artifacts within large-scale image-text datasets that can trigger misclassification.
To that end, we propose the search pipeline in Sec. 3.1, analyze the extracted artifacts to reveal broader insights into model vulnerabilities in Sec. 3.2, introduce an evaluation metric that accounts for occlusion effects in Sec. 3.3, and finally, propose a mitigation strategy in Sec. 5.
Figure 3: Examples of Attack Artifacts split into: graphics without text ($A^{\text{graphics-no-text}}_{c^{*}}$), graphics with embedded text ($A^{\text{graphics-text}}_{c^{*}}$), and unrelated text ($A^{\text{text}}_{c^{*}}$). Each row corresponds to a different target class (Brown Hair, Japan, Adult), that models have learned to associate with the artifacts. Notably, text artifacts need not match the class exactly, while graphical symbols can represent indirect but learned associations. These findings highlight the diverse range of artifacts that can manipulate model predictions. Refer to Supplementary. A for more examples.
@@ 3.1 Attack Pipeline
To systematically search for attack artifacts, we develop a search pipeline. First, we extract artifacts from pretraining datasets by filtering images containing either text or graphical elements, categorizing them into non-matching text, graphics with text, and graphics without text (Sec. 3.1.1). Next, we evaluate how these artifacts influence model predictions by applying them to opposing-class images and selecting those that consistently increase misclassification rates (Sec. 3.1.2). Finally, we optimize artifact placement to maximize attack effectiveness (Sec. 3.1.3).
@@ 3.1.1 Extracting Artifacts from a Pretraining Dataset
Next, we categorize artifacts into three types, as shown in Fig. 2 Green : unrelated text, graphics with embedded text, and graphics without text. To achieve this categorization, we apply OCR-based text detection to identify artifacts containing text and separate them from purely graphical artifacts. Additionally, to ensure that the extracted text does not merely replicate typographic attacks from prior work [2, 25], we filter out any text that explicitly matches class names in the downstream dataset. This structured approach allows us to systematically evaluate how VLMs respond to textual and graphical cues, both independently and in combination, providing deeper insight into their learned vulnerabilities.
@@ 3.1.2 Finding Artifacts That Influence a Target Class
After obtaining our set of artifacts in Sec. 3.1.1, the next step is to determine which ones are most effective at misleading model predictions. Not all artifacts significantly influence classification—some may be ignored by the model, while others may reinforce correct predictions. To systematically identify high-impact artifacts, we evaluate their ability to shift model predictions toward a target class.
To that end, we apply each artifact to opposing-class images (e.g. if the target class is "car images" then we use every image that is not a car.) and measure its effectiveness in shifting predictions toward the target class as Fig. 2 Purple illustrates. If an artifact consistently causes images from unrelated classes to be classified as the target class, it reveals a strong learned correlation in the model that can be exploited for adversarial purposes [24]. Formally, given a downstream dataset $\mathcal{D}_{\text{downstream}}$, we define a subset of opposing-class images as:
@@
where higher values indicate a stronger influence on misclassification. We compute this score for each artifact in our three categories: text artifacts ($\mathcal{A}^{\text{text}}$), graphics with embedded text ($\mathcal{A}^{\text{graphics-text}}$), and graphics without text ($\mathcal{A}^{\text{graphics-no-text}}$), as defined in Sec. 3.1.1. For each artifact category, we rank artifacts by $P_{c^{*}}(a)$ in descending order and select the top $s$ most effective ones. This results in the sets $\mathcal{A}_{c^{*}}^{\text{text}}$, $\mathcal{A}_{c^{*}}^{\text{graphics-text}}$, and $\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$, where $c^{*}$ denotes the target class.
@@ 3.1.3 Optimizing Artifact Placement
While Sec. 3.1.2 identified artifacts that are consistently effective across different placements, their positioning within an image can further enhance misclassification rates.
To address this, for each artifact $a\in\mathcal{A}_{c^{*}}^{\text{top-text}}\cup\mathcal{A}_{c^{*}}^{\text{top-graphics-text}}\cup\mathcal{A}_{c^{*}}^{\text{top-graphics-no-text}}$, we evaluate a set of candidate locations $\mathcal{L}$ to determine the most effective positioning. For each artifact, we pick the location that obtains the highest success rate. This step fine-tunes the effect of already strong artifacts, maximizing misclassification rates. Refer to the Supplementary B for an ablation of location effect. In short, we find that the top border of the image (especially top middle) to be the most effective.
@@ 3.2 Insights from the Artifacts
After extracting and ranking artifacts based on their influence on model predictions, we examine the highest-impact artifacts for different target classes. Fig. 3 presents examples of these artifacts, categorized into graphics without text ($\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$), graphics with embedded text ($\mathcal{A}_{c^{*}}^{\text{graphics-text}}$), and unrelated text artifacts ($\mathcal{A}_{c^{*}}^{\text{text}}$). These examples reveal distinct patterns in how different artifact types induce misclassification.
Non-textual graphics often share visual attributes loosely related to the target class. For instance, brown-colored logos and animal symbols appear frequently for "Brown Hair," while "Japan" is associated with red-and-white designs reminiscent of its national flag. Similarly, the "Senior" class is linked to medical symbols and retirement-related imagery, suggesting that models rely on broad visual associations learned from pretraining data.
Graphics with embedded text frequently contain contextually related words, even when the complete phrase is irrelevant. Examples include “Brown” from unrelated brand names appearing in the "Brown Hair" category and “Japan Cloud” for the "Japan" category. This demonstrates that models are highly sensitive to the presence of text fragments within graphical elements, further reinforcing the text-heavy biases introduced during pretraining.

### v2

@@ Abstract
Vision-language models (VLMs) (e.g., CLIP, LLaVA) are trained on large-scale, lightly curated web datasets, leading them to learn unintended correlations between semantic concepts and unrelated visual signals. These associations degrade model accuracy by causing predictions to rely on incidental patterns rather than genuine visual understanding. Prior work has weaponized these correlations as an attack vector to manipulate model predictions, such as inserting a deceiving class text onto the image in a “typographic” attack. These attacks succeed due to VLMs’ text-heavy bias—a result of captions that echo visible words rather than describing content. However, this attack has focused solely on text that matches the target class exactly, overlooking a broader range of correlations, including non-matching text and graphical symbols, which arise from the abundance of branding content in web-scale data. To address this gap, we introduce “artifact-based” attacks: a novel class of manipulations that mislead models using both non-matching text and graphical elements. Unlike typographic attacks, these artifacts are not predefined, making them simultaneously harder to defend against and more challenging to find. We address this by framing artifact attacks as a search problem and demonstrate their effectiveness across five datasets, with some artifacts reinforcing each other to reach 100% attack success rates. These attacks transfer across models with up to 90% effectiveness, making it possible to attack unseen models. To defend against these attacks, we extend prior work’s artifact aware prompting to the graphical setting. We see a moderate reduction of success rates of up to 15% relative to standard prompts, suggesting a promising direction for enhancing model robustness. Code: https://github.com/mqraitem/Web-Artifact-Attacks
@@ 1 Introduction
Figure 1: We propose a novel class of Web Artifact Attacks that expand the attack surface of Vision-Language Models beyond prior typographic attacks. For a sample task of predicting a person’s age, (a) illustrates Typographic Attacks from prior work [3] which manipulate predictions using class-matching text. In contrast, our proposed web artifact attacks, consisting of (b) non-matching text, (c) standalone graphics, and (d) graphics with embedded text, can also mislead the model.
Figure 2: Overview of the Web Artifacts Search and Selection. The process involves two steps: Step 1: We filter an image-text dataset using CLIP-based retrieval and a set of artifact prompts. Retrieved artifacts are categorized into three groups: graphics with text ($\mathcal{A}^{\text{graphics-text}}$), graphics without text ($\mathcal{A}^{\text{graphics-no-text}}$), and unrelated text ($\mathcal{A}^{\text{text}}$) using OCR-based text extraction (Sec.˜3.1.1). Step 2: artifacts are applied to opposing-class images (e.g., "Smiling" images when targeting "Frowning"), and their effectiveness is evaluated based on their ability to mislead model predictions. Artifacts are ranked according to their attack success rate, resulting in a refined set of high-impact artifacts for each category ($\mathcal{A}_{c^{*}}^{\text{graphics-text}}$, $\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$, and $\mathcal{A}_{c^{*}}^{\text{text}}$) (Sec.˜3.1.2).
@@
Vision-language models (VLMs), such as CLIP [28] and LLaVA [19], are trained on large-scale, lightly curated web datasets [31, 33]. Since web captions often describe what is present rather than what is important, models learn spurious patterns—unintended correlations that arise from frequent but unrelated co-occurrences [38, 1, 12, 14, 35, 34, 23, 16, 18]. For example, prior work found that Chinese characters frequently co-occur with the word “carton” in ImageNet [16], leading the model to misclassify images containing Chinese characters as a “carton,” regardless of its actual content. Moreover, these correlations can be weaponized by attackers to manipulate model predictions. Fig.˜1(a) shows a well-documented example, typographic attacks [27, 3], where inserting the text of the wrong category into an image tricks the model into predicting that category. This attack takes advantage of the model’s over-reliance on text, a bias reinforced by pretraining datasets in which captions frequently parrot text visible in the images [18].
@@
A shortcoming of prior work is that they focused only on explicit class-matching text, overlooking a broader set of vulnerabilities, including non-matching text and graphical artifacts such as logos. These elements, which are common in web-scale datasets due to branding and advertisements, may introduce unintended visual cues that models may rely on for classification. To address this gap, in Fig.˜1, we introduce Web Artifact Attacks, a broader class of attacks that exploit both non-matching text (Fig.˜1(b)) and graphics (Fig.˜1(c) and (d)) to mislead VLM predictions. We categorize graphical artifacts into two types: those with embedded text, which may reinforce textual biases, and those without text, isolating the impact of purely visual cues.
@@
Unlike typographic attacks [27, 3], where the adversary inserts class-matching text to deliberately exploit the model’s reliance on words, Web Artifact Attacks pose a greater threat because the defender does not know in advance the attacking artifacts. Thus, it is not obvious how to identify effective attacking artifacts or defend against them. To find attacks, we frame Web Artifact Attacks as a search problem, scanning image-text datasets [30, 6] to uncover text and graphics that can distort model predictions. The simplicity of these attacks makes them a notable security concern for deployed VLMs, as they only require dataset queries using similarity retrieval, making them highly accessible.
@@
To that end, our search process consists of three stages: (1) Fig.˜2 Green: Artifact Retrieval, where we extract candidate artifacts by scanning image-text datasets for images with text or graphics (2) Fig.˜2 Purple: Effect Estimation, where we apply artifacts to images from opposing classes and evaluate their influence on model predictions, and (3) placement optimization, where we refine the positioning of artifacts to maximize attack success.
@@
We evaluate the resulting attack artifacts across different VLM training objectives, visual encoder architectures, and pretraining dataset curation. Our findings demonstrate that Web Artifact Attacks are highly effective across various VLM training paradigms, including contrastive learning [28], sigmoid loss [37], and Large Language Model (LLM) fusion with visual features [19, 39]. Furthermore, these vulnerabilities persist regardless of the visual encoder size (ViT-B-32, ViT-B-16, ViT-L-14), and attempts at stronger pretraining dataset curation (DataComp [11]).
@@
Finally, we explore mitigation strategies by building on prior work [8], which showed that incorporating typographic attacks into the prompt can reduce their effectiveness. For example, if an image of a youth is attacked with the text “senior,” the prompt “an image of a senior with the word ‘youth’ written on it” can weaken the attack. However, unlike typographic attacks, our attacks include graphical elements that cannot be easily incorporated into text. To address this, we propose augmenting the prompt with descriptions of graphical artifacts obtained through a captioning model. For LLM-based VLMs, we prompt the model to identify and describe these artifacts in its response. This approach reduces success rates by up to 15% relative to standard prompts, suggesting a promising direction for enhancing model robustness.
@@
  • We expose Web Artifact Attacks, which extend beyond typographic attacks by leveraging non-matching text, standalone graphics, and graphics with embedded text to mislead VLM predictions.
@@ 2 Related Work
Relation to Adversarial Attacks. Adversarial attacks [36, 29, 40, 10, 7] learn an imperceptible noise, that disrupts the model visual recognition capabilities when added to the model. More recently, adversarial techniques have been weaponized for jailbreaking multimodal models to bypass safety mechanisms and generating restricted outputs [25, 32, 4]. While our attacks share a similar end goal to adversarial attacks, namely, disrupting model behavior, adversarial and artifact attacks are fundamentally different from each other. Our attacks originate as a result of spurious correlation in uncurated pretraining datasets, while adversarial attacks result from an artificial noise uniquely designed through the use of model intrinsics (e.g., gradients) to disrupt the fundamental mechanics of neural nets. Therefore, to develop such attacks, an attacker needs a sophisticated understanding of the intrinsic mechanics of neural nets such as e.g., Gradient [36, 29, 40], Meta Learning [10] or Attention maps [7] to name a few. However, the attacker, in our setup, only needs to know how to query the model.
@@ 3 Web Artifact Attacks Against Vision-Language Models
We propose Web Artifact Attacks, a new vector of unpredictable class of attacks that leverage artifacts from image-text datasets [30, 6]. Unlike typographic attacks, these artifacts—such as non-matching text (text that contains partial overlaps, phonetic similarities, or visually similar characters but does not explicitly spell out the target class), graphical symbols—do not need to explicitly match the target class yet can still mislead model predictions.
@@
Since these artifacts are not known a priori, they are harder to defend against while also making it challenging to determine in advance which artifacts will be effective. Therefore, we frame finding our attack artifacts as a search problem: systematically identifying artifacts within large-scale image-text datasets that can trigger misclassification.
To that end, Sec.˜3.1 describes our proposed the search pipeline, Sec.˜3.2 analyzes the extracted artifacts to reveal broader insights into model vulnerabilities, Sec.˜3.3 introduces an evaluation metric that accounts for occlusion effects, and finally, Sec.˜5 proposes a mitigation strategy.
@@ 3.1 Attack Pipeline
This section describes our search pipeline to systematically find attack artifacts. First, in Sec.˜3.1.1, we extract artifacts from pretraining datasets by filtering images containing either text or graphical elements, categorizing them into non-matching text, graphics with text, and graphics without text. Next, in Sec.˜3.1.2, we evaluate how these artifacts influence model predictions by applying them to opposing-class images and selecting those that consistently increase misclassification rates. Finally, in Sec.˜3.1.3, we optimize artifact placement to maximize attack effectiveness.
@@ 3.1.1 Extracting Artifacts from a Pretraining Dataset
Next, as shown in Fig.˜2, we categorize artifacts into three types Green : unrelated text, graphics with embedded text, and graphics without text. To achieve this categorization, we apply OCR-based text detection (We use off-the-shelf models from EasyOCR [2]) to identify artifacts containing text and separate them from purely graphical artifacts. Additionally, to ensure that the extracted text does not merely replicate typographic attacks from prior work [3, 27], we filter out any text that explicitly matches class names in the downstream dataset. This structured approach allows us to systematically evaluate how VLMs respond to textual and graphical cues, both independently and in combination, providing deeper insight into their learned vulnerabilities.
@@ 3.1.2 Finding Artifacts That Influence a Target Class
After obtaining our set of artifacts in Sec.˜3.1.1, the next step is to determine which ones are most effective at misleading model predictions. Not all artifacts significantly influence classification—some may be ignored by the model, while others may reinforce correct predictions. To systematically identify high-impact artifacts, we evaluate their ability to shift model predictions toward a target class.
To that end, as Fig.˜2 Purple illustrates, we apply each artifact to opposing-class images (e.g. if the target class is “car images” then we use every image that is not a car) and measure its effectiveness in shifting predictions toward the target class If an artifact consistently causes images from unrelated classes to be classified as the target class, it reveals a strong learned correlation in the model that can be exploited for adversarial purposes [26]. Formally, given a downstream dataset $\mathcal{D}_{\text{downstream}}$, we define a subset of opposing-class images as:
@@
where higher values indicate a stronger influence on misclassification. We compute this score for each artifact in our three categories: text artifacts ($\mathcal{A}^{\text{text}}$), graphics with embedded text ($\mathcal{A}^{\text{graphics-text}}$), and graphics without text ($\mathcal{A}^{\text{graphics-no-text}}$), as defined in Sec.˜3.1.1. For each artifact category, we rank artifacts by $P_{c^{*}}(a)$ in descending order and select the top $s$ most effective ones. This results in the sets $\mathcal{A}_{c^{*}}^{\text{text}}$, $\mathcal{A}_{c^{*}}^{\text{graphics-text}}$, and $\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$, where $c^{*}$ denotes the target class.
@@ 3.1.3 Optimizing Artifact Placement
While Sec.˜3.1.2 identified artifacts that are consistently effective across different placements, their positioning within an image can further enhance misclassification rates.
To address this, for each artifact $a\in\mathcal{A}_{c^{*}}^{\text{top-text}}\cup\mathcal{A}_{c^{*}}^{\text{top-graphics-text}}\cup\mathcal{A}_{c^{*}}^{\text{top-graphics-no-text}}$, we evaluate a set of candidate locations $\mathcal{L}$ to determine the most effective positioning. For each artifact, we pick the location that obtains the highest success rate. This step fine-tunes the effect of already strong artifacts, maximizing misclassification rates. Refer to the Supp. B for an ablation of location effect. In short, we find that the top border of the image (especially top middle) to be the most effective.
@@ 3.2 Insights from the Artifacts
Figure 3: Examples of Web Artifacts split into: graphics without text ($A^{\text{graphics-no-text}}_{c^{*}}$), graphics with embedded text ($A^{\text{graphics-text}}_{c^{*}}$), and unrelated text ($A^{\text{text}}_{c^{*}}$). Each row corresponds to a different target class (Brown Hair, Japan, Adult), that models have learned to associate with the artifacts. Notably, text artifacts need not match the class exactly, while graphical symbols can represent indirect but learned associations. These findings highlight the diverse range of artifacts that can manipulate model predictions. Refer to Supp. A for more examples.
After extracting and ranking artifacts based on their influence on model predictions, we examine the highest-impact artifacts for different target classes. Fig.˜3 presents examples of these artifacts, categorized into graphics without text ($\mathcal{A}_{c^{*}}^{\text{graphics-no-text}}$), graphics with embedded text ($\mathcal{A}_{c^{*}}^{\text{graphics-text}}$), and unrelated text artifacts ($\mathcal{A}_{c^{*}}^{\text{text}}$). These examples reveal distinct patterns in how different artifact types induce misclassification.
Non-textual graphics often share visual attributes loosely related to the target class. For instance, brown-colored logos and animal symbols appear frequently for “Brown Hair,” while “Japan” is associated with red-and-white designs reminiscent of its national flag. Similarly, the “Senior” class is linked to medical symbols and retirement-related imagery, suggesting that models rely on broad visual associations learned from pretraining data.
Graphics with embedded text frequently contain contextually related words, even when the complete phrase is irrelevant. Examples include “Brown” from unrelated brand names appearing in the “Brown Hair” category and “Japan Cloud” for the “Japan” category. This demonstrates that models are highly sensitive to the presence of text fragments within graphical elements, further reinforcing the text-heavy biases introduced during pretraining.